find the smallest whole number which when divided by 6 gives a remainder of 1 and when divided by 11 gives a remainder of 6.
Answer by ankor@dixie-net.com(22740)

find the smallest whole number which when divided by 6 gives a remainder of 1
and when divided by 11 gives a remainder of 6.
:
One way is to make list of the numbers as follow
Write the multiples of 6 plus 1; first one 24+1 = 25, 30+1 = 31 etc
25 + 31 + 37 + 43 + 49 + 55 + 61 + 67 + 73
:
Write the multiples of 11 plus 6; ie 22+6 = 28, 33+6 = 39 etc
28, 39, 50, 61, 72, 83, 94
:
Look for the first number that occurs in both lists
61 is our number